Your favourite chocolate bar contains 1110 kJ of energy. If you go for a vigorous bike ride your metabolic rate will increase to 400 W. How long will you have to ride to use up the 1110 kJ from the chocolate bar (in minutes)?


1
Expert's answer
2021-02-02T17:05:42-0500

By the definition of the power, we have:


"P=\dfrac{W}{t},""t=\dfrac{1110\cdot10^3\ J}{400\ W}=2775\ s\cdot{1\ min}{60\ s}=46\ min."